OpenStudy (anonymous):

What is the median of all the numbers 5440,2533,2685,2893,4522,2320,3597,2631,2222, 3665,5255,3494 Remember when you wanna find the median your data should be set from least to greatest. :)

OpenStudy (anonymous):

\[{2222, 2320, 2533, 2631, 2685, 2893, 3494, 3597, 3665, 4522, 5255, 5440}\]

OpenStudy (anonymous):

Correct, now just start to cross out

OpenStudy (anonymous):

Even amount of numbers, so you have to find two numbers

OpenStudy (anonymous):

is the median 2893

OpenStudy (anonymous):

\[\frac{2893+3494}{2}=\frac{6387}{2}=3193.5 \] http://www.mathsisfun.com/median.html

OpenStudy (anonymous):

so the median is 3193.5